OpenStudy (anonymous):

how do I convert r^2=2sin2theta to Cartesian form

OpenStudy (anonymous):

r^2 = 2sin(2theta) r^2 = 2 * 2sin(theta) cos(theta) r^2 = x^2 + y^2 cos(theta) = x/r sin(theta) = y/r (x^2+y^2)^2 = 4 (y/r) (x/r) (x^2 + y^2)^2 = 4xy / r^2 r^2 (x^2 + y^2)^2 = 4xy (x^2 + y^2)^2 (x^2 + y^2)^2 = 4xy (x^2 + y^2)^2

OpenStudy (anonymous):

(x^2 + y^2)^4 = 4xy